OpenStudy (anonymous):

What is the value of x? 3x^3 − 6=186

OpenStudy (campbell_st):

well start by solving for x^3 3x^3 - 6 = 186 _ 6 +6 -------------- 3x^3 = 192 divide both sides by 3 then take the cube root

OpenStudy (anonymous):

i got 64 but thats not the answer i know that

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

ok.. so you have \[x^3 = 64\] now take the cube root of both sides

OpenStudy (anonymous):

how do i do that @campbell_st

OpenStudy (campbell_st):

well you need \[x = \sqrt[3]{64}\]

OpenStudy (anonymous):

is it 4 @campbell_st

OpenStudy (campbell_st):

it is.... 4 x 4 x 4 = 64

OpenStudy (anonymous):

okay thanks